OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2ð). cos(4x) - cos(2x) = 0

OpenStudy (luigi0210):

is that an exponent or no?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@Luigi0210 @dumbcow can u help me

OpenStudy (dumbcow):

use double angle formula \[\cos 2 \theta = 2\cos^{2} \theta -1\] \[\rightarrow (2\cos^{2} (2x) -1)-\cos(2x) = 0\] treat it like a quadratic equation and solve for "cos 2x "

OpenStudy (anonymous):

How do u write it as a quad

OpenStudy (dumbcow):

\[2u^{2} -u-1 = 0\] where u=cos 2x this will factor

OpenStudy (anonymous):

how do u get cos2x = u

OpenStudy (dumbcow):

magic ...haha no i made it up to help you see the quadratic form its called a substitution

OpenStudy (anonymous):

so i dont need the cos2x

OpenStudy (anonymous):

am i just subbing it because i can

OpenStudy (dumbcow):

exactly, it helps to factor it ... but in the end you will have ...cos 2x = ?

OpenStudy (anonymous):

now i have 2x=0 2x=cos of (-1/2)

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@torobi From where @dumbcow left off, after you have factored the quadratic where you made the substitution u = cos(2x), you get:\[\bf 2u^2-u-1=0 \implies (2u+1)(u-1)=0 \implies u = -\frac{1}{2} \ or \ u=1\]After we have factored and achieved our solution, we bring cos(2x) back in and replace u with cos(2x) again:\[\bf \cos(2x)=u \implies \cos(2x)=-\frac{1}{2}\]\[\bf \cos(2x)=u \implies \cos(2x)=1\]Now we solve for 'x' in each case and get our solutions to whatever 'x' is. Can you do that? @torobi